OpenStudy (anonymous):

What are the major steps to solving a related rates problem in Calc 1?

OpenStudy (anonymous):

finding an initial equation that contains the quantities that will be undergoing change. ie if asked to find how fast the volume of a sphere is changing as the radius changes, you will need an intitial equation containing both Volume and Radius

OpenStudy (anonymous):

Taking implicit differentiation wrt to time is often required due to the rate, a time component

OpenStudy (anonymous):

1) before even finding an equation you should draw the picture... this sometimes helps to figure out what you know and what you're trying to find. For example if you're trying to find the rate a ladder's legs are slipping from the wall and you're given that the ladder lays out to height y

OpenStudy (anonymous):



OpenStudy (anonymous):

then you use what you know to formulate an equation ... sometimes you can put stuff in before implicitly deriving.. if that value is fixed (such as the ladder not growing ), you can put in into the equation. for example the equation i'd use for this is \[a^2+b^2=c^2\] since hypotenuse doesn't get bigger or smaller, it can be put into the equation right away \[a^2+b^2=k\] where k is a constant now when you implicitly solve you'll get 0 on the right side..

OpenStudy (anonymous):

Thank you guys! ;D